A parallel-plate capacitor is made from two aluminum-foil sheets, each 6.7 cm wide and 5.5 m long. Between the sheets is a Teflon strip of the same width and length that is 4.5×10−2 mmthick.

What is the capacitance of this capacitor? (The dielectric constant of Teflon is 2.1.)
Answer #1

A=0.067*5.5=0.368m^2

d=4.5×10−5 m

c =epison *dielectric constant*A/d
c= 8.89*10^-12 *2.1 * 0.368 /4*10^-5

=1.717*10^-7 F

=0.1717uF